if the sum of first n terms of an A.P. is cn2 ,then the sum of squares of these n terms is 
(A) n(4n2 -1)c2/2
(B) n(4n2 +1)c2/3
(C) n(4n2 -1)c2/3
(D) n(4n2 +1)c2/6
 

Grade:12

2 Answers

Riddhish Bhalodia
8 years ago
A be the first term and d be the difference of the AP
sum of n terms is
n(A + d(n-1)/2) = cn2
simplifying we get the equation as
(A-d/2) + nd/2 = cn
now comparing coefficients in ‘n’ we get two relations as
A - d/2 = 0 and d/2 = c
thus A = c and d = 2c
then the AP (in terms of c and n) is as follows
c , 3c , 5c , 7c , ….. , (2n-1)c
Now we need to evaluate the sum of squares of these n terms i.e as
S = (12 + 32 + …. + (2n-1)2) c2
using the formula for the sum of squares of first n odd numbers we get the final value of S as
S = \frac{c^2(2n-1)(2n+1)n}{3} = \frac{c^2n (4n^2-1)}{3}
Hence the option C is correct answer
Prabhakar ch
8 years ago
The answer for your question is C bye all the best for your bright future................................................